OpenStudy (anonymous):

ABCD is a parallelogram AB= 2x+5, BC=3x+7 , CD= 3x-8 , how long are AD and AB?

OpenStudy (anonymous):



OpenStudy (anonymous):

Since it is a parallelogram, sides CD and AB are of equal length... so set your equation up 2x+5 = 3x - 8 and solve for x.

OpenStudy (anonymous):

thankyouuuuuuuuu :*

OpenStudy (anonymous):

are you able to solve for x?

OpenStudy (anonymous):

i'm trying? but i think i still can't.

OpenStudy (anonymous):

well since it is -8, then you have to +8 on the other side... so 5+8 = 3x - 2x

OpenStudy (anonymous):

so you are left with 13 = 1x... and then you are just left to substitute back into each expression to get your side lengths.

OpenStudy (anonymous):

yah yah, i forgot , sorry thanks again :**

OpenStudy (anonymous):

yw... anytime.

OpenStudy (anonymous):

last question puhleaseeee? ? abcd is a parallelogram if angle a=3x-10, and angle c=2x+98 find x , angle a & angle b.

OpenStudy (anonymous):

well the angles of a parallelogram are either going to be (1) supplementary or (2) congruent, just looking at my diagram, do you know the relationship between angle a and angle c, are those congruent or supplementary?

OpenStudy (anonymous):

congruent. so 3x+10=2x+98 ,and the x=88 am i ryt? :)

OpenStudy (anonymous):

well right above that you said that it was 3x - 10, but yes you set them equal to each other and solve for x....

OpenStudy (mathlegend):

@zonazoo when you said plug back in the 13 for x. It said find side AD. So would that side just be 13? Or would I move the..... 3x+7 over there since it is parallel?

OpenStudy (anonymous):

yes, when you plug that into 3(13)+7 for BC, that value will be the value of AD since those two sides are congruent.... so AD = 46.

OpenStudy (anonymous):

thankyou :*

OpenStudy (anonymous):

yw